### v1

# Unveiling the Sagittarius Dwarf Spheroidal Galaxy Core with Gaia DR3

Abstract
The Sagittarius dwarf spheroidal galaxy provides us with the unique opportunity to study an ongoing Galactic cannibalistic event between our Milky Way Galaxy and a satellite dwarf galaxy. We present the first major membership star catalog of the Sagittarius dwarf core (400,000 sources) and Messier 54 (2000 sources) with positions, proper motions, and parallaxes from DR3, supplemented with metallicities from the Apache Point Observatory Galactic Evolution Experiment. We isolate the Sagittarius dwarf core (3.7 core radii) and Messier 54 (0.125 core radii) using literature positions. Using evolutionary sub-samples separated within a color-magnitude diagram, we analyze the structures and substructures of the Sagittarius core and infer its positional relationship with Messier 54 within 5D phase space. A sample of Milky Way stars from a similar galactic latitude was used to identify contaminants and separate member stars from the core of the Sgr dSph and Messier 54 using a Gaussian distribution in proper motion space. We present the derived proper motion, parallaxes, and metallicities for these evolutionary sub-samples while demonstrating the precision of our sample using red clump standard candles. We find a distance modulus for the Sagittarius core and Messier 54 of mag and mag, corresponding to a heliocentric distance of kpc and kpc respectively. With red clump distance analysis, our results imply there is no separation between the Sagittarius core and Messier 54.
1 Introduction
The Sagittarius dwarf spheroidal galaxy (Sgr dSph, hereafter Sgr, Ibata et al., 1994, 1995), discovered nearly 30 years ago, is the closest known example of a dwarf satellite galaxy undergoing tidal disruption and accretion into a larger galaxy–the Milky Way (MW) (Johnston et al., 1995; Zhao, 1998; Johnston, 1998; Helmi et al., 1999; Bellazzini et al., 2003; Putman et al., 2004; Lewis & Ibata, 2005; Bellazzini et al., 2008; Laporte et al., 2018; Bellazzini et al., 2020; Horta et al., 2023). Dwarf galaxies are considered the building blocks of the hierarchical merging process (Gallagher & Wyse, 1994; Tosi, 2003) and, thus, are fundamental mechanisms for forming large galaxies through a process of galactic cannibalism for spiral galaxies (Hausman & Ostriker, 1978; Helmi, 2001). Tidal interactions during close encounters with large host galaxies often transform star-forming, gas-rich dwarf irregular (dIrr) galaxies with small rotating disks into classical, gas-depleted, pressure-supported dwarf spheroidals (Frinchaboy et al., 2012). In the case of Sgr, these forces from the MW’s tidal field have elongated its core toward the Galactic plane (Majewski et al., 2003), complicating observations due to its proximity of kpc below (Mucciarelli et al., 2017) and kpc behind the MW bulge (Ibata, 1999). Three pericentric passages with the MW have stripped the Sgr system, creating two massive tidal streams—a trailing arm in the south-galactic hemisphere and a leading arm in the north-galactic hemisphere–populating the MW halo with 5% of its halo M giant stars and associated dark matter particles (Majewski et al., 2003; Bellazzini et al., 2020).
These pericentric passages and crossings through the MW disk have coincided with increased periods of star formation within Sgr and the MW, occurring 5.9, 1.9, and 1 Gyr ago (de Boer et al., 2015; Ruiz-Lara et al., 2020). These episodes have led to a dominance of intermediate-age stars within the Sgr core, aged around 4 to 6 Gyr, with metallicities ranging from [Fe/H] = to (Siegel et al., 2007). A younger population with [Fe/H] = formed approximately 1 Gyr ago, aligns with the most recent pericentric passage, and metal-poor old population of [Fe/H] = (Siegel et al., 2007; Alfaro-Cuello et al., 2019). Although material from Sgr is dispersed across a wide range of distances (10-100 kpc) due to the tidal disruption, parts of the main body remain embedded within these tidal streams, carrying stars and dark matter lost during this process (Hasselquist et al., 2017; Fernández-Trincado et al., 2021). While understanding the tidal streams of Sgr are crucial for probing the MW gravitational potentials and changes within the disk, understanding the core dynamics, physical properties, and kinematics, of the Sgr main body is essential for deciphering the evolution and nature of the Sgr system. Investigating these properties through galactic archaeology—studying galactic structure and evolution by examining the ages, chemical compositions, and distances of stellar populations—offers insight into Sgr and its relationship with M54.
The main body (including the core) of Sgr also exhibits distinctive features. According to del Pino et al. (2021), Sgr has a bar approximately 2.5 kpc in length, inclined at in the plane of the sky, with tidal tails extending in an ”S” shape from the ends of the bar. Despite significant tidal stripping, Sgr retains a minor clockwise rotation, which extends along the outer regions of its tidal stream (del Pino et al., 2021). Currently, four globular clusters are believed to be within the main body of Sgr: Messier 54 (M54), Arp 2, Terzan 7 (Ter 7), and Terzan 8 (Ter 8) (Bellazzini et al., 2020). M54 (NGC 6715) lies at the photometric center of the Sgr core and has been proposed as the nucleated core of Sgr due to also lying within the densest region of the Sgr core (Monaco et al., 2005). However, further studies (Bellazzini et al., 2008; Kunder & Chaboyer, 2009; del Pino et al., 2021; An et al., 2024) suggest that M54 formed independently of Sgr’s nucleus, potentially being captured during tidal disruption, with Siegel et al. (2011) inferring a kpc separation between the Sgr core and M54. Supporting this, Minelli et al. (2023) found that M54’s mean metallicity ([Fe/H] ), differing by approximately one full dex from the median metallicity of the Sgr core found in Hayes et al. (2020) of [Fe/H] = , reinforcing the idea of M54 forming independently of Sgr. Meanwhile, the central region of Sgr reveals a complex formation history, displaying a mix of young ( Gyr), intermediate-age (around 4–6 Gyr), and old ( Gyr) stellar populations, with metallicities spanning from [Fe/H] (Siegel et al., 2007; Alfaro-Cuello et al., 2019; An et al., 2024).
Following the release of DR3, Gaia Collaboration et al. (2021) probed the structure of the LMC and SMC, showcasing improvements from Gaia Collaboration et al. (2018) where they previously determined the kinematics, including mean proper motion (, ) and parallax of 75 Galactic globular clusters, the Large and Small Magellanic clouds (LMC and SMC respectively), and nine dwarf spheroidal galaxies including the Sgr core within DR2. They found that photometry formerly affected by background issues caused by high stellar density in central areas of dwarfs were reduced, there was a two-fold reduction in proper motion uncertainty, and an overall increase in the precision of astrometry and photometry with significantly reduced systematic effects from DR2. APOGEE DR17 (Abdurro’uf et al., 2022a) contains measurements for stars from dwarf satellite galaxies, open clusters, and globular clusters (GCs) identified from procedures detailed in Zasowski et al. (2013), including Sgr and M54. The mixture of these two sky surveys allows us to constrain Sgr and M54 member stars from the MW disk to analyze the dynamics within 5D phase space.
In this work, we describe membership selection of the Sgr core and M54 using DR3 and APOGEE DR17, inspired by the methodology within Gaia Collaboration et al. (2018, 2021). We describe the data selection procedure in Section 2. In Section 3 we describe the spatial separation process, removing background MW contaminants, and isolating evolutionary areas of the Sgr core and M54. In Section 4, we use our sample of red clump (RC) stars to estimate a distance to Sgr and M54 and present our result from our distance determination. In Section 5, we present results from 5D phase space analysis of the Sgr core and M54, compare our distance determinations from the RC with prior studies. We compare our membership sample to previous work and discuss the efficacy of our sample for future studies. In Section 6, we summarize our conclusions.
2 Data
2.1 Gaia DR3
We use data from the third data release ( DR3; Gaia Collaboration et al. (2016, 2023)), for the five-parameter astrometric solution; Creevey et al. (2023); Gaia Collaboration et al. (2023); right ascension (), declination (), parallax (), proper motion in right ascension direction (; defined as ), and proper motion in declination direction ().
We extract a sample from the DR3 catalog based on specific spatial criteria, where , are the accepted center of the Sgr core from literature in right ascension and declination respectively using the coordinates () = (283.75∘, -30.46∘) (Majewski et al., 2013), with a radius of . This initial large selection radius is greater than defined in Majewski et al. (2013) as the core of Sgr has a defined ellipticity of (Majewski et al., 2013). We constrain our selection to sources that have at least a five-parameter astrometric solution. This selection can be reproduced in full using the ADQL query within Appendix 8.1. The resulting sample contains 4,378,343 objects.
2.2 SDSS/APOGEE DR17
We also use data from the Apache Point Observatory Galactic Evolution Experiment (APOGEE 1 & 2; Majewski et al., 2013), under the Sloan Digital Sky Survey III & IV (Eisenstein et al., 2011; Blanton et al., 2017). Observations were conducted using the 2.5 m Sloan Foundation Telescope (Gunn et al., 2006) at Apache Point Observatory and the 2.5 m du Pont telescope (Bowen & Vaughan, 1973) at Las Campanas Observatory, with data collected by the APOGEE-North and APOGEE-South spectrographs (Wilson et al., 2019). The selection process for APOGEE targets is described in Zasowski et al. (2013, 2017); Beaton et al. (2021); Santana et al. (2021).
The APOGEE spectra were processed using the APOGEE reduction pipeline (Nidever et al., 2015), and key parameters were derived through the APOGEE Stellar Parameters and Chemical Abundances Pipeline (ASPCAP; Garc´ıa Pérez et al., 2016). Further details on the spectral analysis process can be found in Shetrone et al. (2015); Smith et al. (2021). The final data release for APOGEE-2 (DR17; Abdurro’uf et al., 2022b) includes approximately 734,000 stars, encompassing all observations from both spectrographs collected between August 2011 and January 2021. A comprehensive overview of the quality and parameter limitations of APOGEE DR17 is available in Abdurro’uf et al. (2022b).
3 Methodology
3.1 Spatial Selection
From our DR3 query, we compute the effective radius of the Sgr core from flattened coordinates defined in Section 8.2, using , the accepted center of the Sgr core from literature (in right-ascension and declination), the ellipticity, and the position angle, as defined from Majewski et al. (2013). This spatial selection includes the globular cluster Messier 54 (M54; NGC6715), which has a tidal radius of 7.5 arcmin (0.125∘) (Trager et al., 1995).
We define our initial sample of Sgr core stars of those within the effective radius of after accounting for the ellipticity of the Sgr core, but outside of the of M54. We defined our initial M54 sample with stars located within of our defined , along with selecting any stars outside of M54’s literature radius if identified within the APOGEE membership flag (MEMBER) (Zasowski et al., 2013, 2017).
To refine our samples and reduce contamination from foreground Milky Way stars, we applied a modified version of the methods described by Gaia Collaboration et al. (2018) and Gaia Collaboration et al. (2021).
1.
To ensure a high-quality five-parameter solution, we excluded stars that did not reach a minimum value of 5 within visibility_periods_used (Lindegren et al., 2018).
2.
We performed a cut in relative parallax error (which is equivalent to ), which corresponds to removing stars within 5 kpc from the Sun. This eliminates stars with high-precision parallax measurements, which are more likely to be foreground objects. DR3 only provides reliable distances out to kpc, literature from Monaco et al., 2004 puts Sgr dSph out to kpc, thus we should keep large parallax errors.
3.
We applied a limiting magnitude cut of G 20.5. This limit is introduced to remove less precise astrometry.
4.
We applied an overall clip to the proper motion in right ascension direction () and declination direction () from the mean PM to further remove any outliers within our sample.
The resulting sample of 1,983,632 objects for Sgr core and 7,852 for M54, while cleaned of foreground sources and potential outliers, still has contaminants from Milky Way background stars. We present the sky density of our initial query sample from after the above quality cuts in the top panel for Sgr dSph and M54 stars in Figure 1.
This contamination is evident in Gaia Collaboration et al. (2018, figure 3) within the color-magnitude diagrams (CMDs) at . We reproduce the figure using our sample defined here, where background objects are also apparent in the top row of CMDs for the Sgr core (left) and M54 (right) within Figure 3.
3.2 Evolutionary Features
Before cleaning of Milky Way contaminants, we isolate evolutionary features in our CMDs for Sgr dSph core and M4. Based on Figure 3 in Gaia Collaboration et al. (2018), we loosely divided our CMDs into three color regions defined as dex, dex dex, and mag (blue, green, and red regions). The blue region appears nearly free of Milky Way contamination, whereas the green and red zones appear heavily contaminated, yet the features of the Sgr and M54 stand out, particularly when viewing stars that have measured [Fe/H] values from APOGEE DR17 in the top panel of Figure 3. We identify the following distinct evolutionary phases within the Sgr core based on Figure 2 from Gaia Collaboration et al. (2021): a young population extending beyond the main sequence turn-off (Young); the blue main sequence (BL MS); the red giant branch (RGB); the asymptotic giant branch (AGB), including long-period variables; the RR Lyrae (RR Lyr) region; the blue horizontal branch (BHB), which also includes RR Lyrae stars; the blue loop (BL), encompassing classical Cepheids; and the red clump (RC). Furthermore, since Sgr is an evolved, disrupted dwarf spheroidal, unlike the Large and Small Magellanic clouds in Gaia Collaboration et al. (2021), we additionally reference the methodology for selecting evolutionary features in de Boer et al. (2012, figure 4) due to the similarity between the Sgr dSph system and the Sculptor dwarf spheroidal system.
Our defined areas are shown in Figure 3, with coordinates for the polygonal areas in Sgr core CMD shown in Table 1 and for M54 within Table 2. There are unassigned regions with the CMD diagrams, as these areas are prevalent of mixing of evolutionary areas, heavy Milky Way contamination, and/or are affected by blended stars. The most notable area of blended stars is within the green and red zones where we predict the main sequence, main sequence turn off, and sub-giant branch would lie, but cannot be disentangled. We note that the CMD for M54 excludes the region for young stars and the blue loop, as these features were not visible. The RGB and AGB within the Sgr core and M54 are noticeable features when viewing the CMDs colored by available [Fe/H] values in Figure 3, where both objects feature distinct RGB and AGB branches of differing metallicities (agreeing with the two branches identified in Minelli et al., 2023, one more metal-poor and one more metal-rich).
| Evolutionary Feature | Polygon Coordinates |
| Young | [0.7, 17.8], [-0.5, 17.8], [0.2, 15], [0.7, 15], [0.7, 17.8] |
| BL MS | [-0.5, 20.5], [-0.5, 19.5], [0.7, 19.5], [0.7, 20.5], [-0.5, 20.5] |
| RGB | [1.2, 19.5], [1.2, 18.1], [1.4, 18.1], [1.4, 17.8], [1.2, 17.8], [1.2, 15], |
| [2.6, 15], [2.0, 16.2], [1.6, 17.8], [3.0, 19.5] | |
| AGB | [1.2, 15], [1.2, 13], [1.5, 13], [4.0, 13], [4.0, 15], [1.2, 15] |
| RR Lyr | [0.4, 17.8], [0.4, 18.4], [0.7, 18.4], [0.7, 17.8], [0.4, 17.8] |
| BHB | [-0.5, 17.8], [-0.5, 18.4], [0.4, 18.4], [0.4, 18.4], [0.4, 17.8], [-0.5, 17.8] |
| BL | [1, 16], [0.7, 15], [0.5, 12], [1.2, 12], [1.2, 13], [1.2, 15] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.4, 17.8], [1.4, 18.1], [1.1, 18.1] |
| Evolutionary Feature | Polygon Coordinates |
| MS | [-0.5, 20.5], [-0.5, 20], [2.5, 20], [2.5, 20.5], [-0.5, 20.5] |
| RGB | [1.2, 19.5], [1.2, 18.1], [1.5, 18.1], [1.5, 17.8], [1.2, 17.8], [1.2, 15], |
| [2.6, 15], [2, 16.2], [1.6, 17.8], [3, 19.5] | |
| AGB | [1.2, 15], [1.2, 14], [1.5, 14], [4.4, 14], [4.4, 15], [1.2, 15] |
| RR Lyr | [0.5, 17.8], [0.5, 18.4], [0.9, 18.4], [0.9, 17.8], [0.5, 17.8] |
| BHB | [-0.5, 17.8], [-0.5, 18.4], [0.5, 18.4], [0.5, 18.4], [0.5, 17.8], [-0.5, 17.8] |
| RC | [1.1, 18.1], [1.1, 17.8], [1.5, 17.8], [1.5, 18.1], [1.1, 18.1] |
| Sgr Core | ||||
| Young | ||||
| BL MS | ||||
| BHB | ||||
| RGB | ||||
| AGB | ||||
| RR Lyr | ||||
| BL | ||||
| RC | ||||
| M54 | ||||
| MS | ||||
| BHB | ||||
| RGB | ||||
| AGB | ||||
| RR Lyr | ||||
| RC |
Note. Parallax is in ; and in ; and in kiloparsecs. The zero-point was corrected for using Lindegren et al., 2021, yet negative parallaxes still occur. Negative parallaxes can be physically interpreted as the source going ’the wrong way around’ on the sky, and are caused by parallaxes with large uncertainties (Luri et al., 2018). For parallax values present here, we adopt a 5- cut to rid our mean values of outliers not fully representative of the main sample.
3.3 Minimizing Milky Way Contamination
The sample of objects obtained in the previous section of the Sgr core and M54 are still heavily contaminated by foreground and background contaminants. To clean our samples of these foreground and background contaminants, we split our larger sample into the identified evolutionary features from our contaminated CMDs in Section 3.2, along with the non-identified evolutionary areas.
From the isolated stars within each evolutionary polygonal area in our CMDs, and all of the stars in unidentified areas we visualize two histograms: one with proper motion in the right ascension direction () and one in proper motion declination direction ().
In each corresponding histogram of and for each defined evolutionary feature and within our identified contaminated areas, there exist two distinct peaks, indicating two distinct populations within our sample as seen in the top and right panel in Figure 2. Using reported literature values from Gaia Collaboration et al. (2018) for the Sgr core within Table C.2, we identify one of these peaks as being Sgr member stars with a and .
To confirm that the other peak present are Milky Way contaminants, we pulled another sample from DR3 at the same galactic latitude of the Sgr core, within a radius, but with an offset of of Galactic longitude, where (referencing and for the Sgr core; Majewski et al. (2013).) We employ the same quality cuts as in Section 3.1, and calculate the respective mean proper motions in the right ascension and declination direction finding, and . These calculated values align with the tallest peak within our histograms as featured in Figure 2.
To remove these Milky Way contaminants, we employ a Gaussian distribution fit taking the peak at the literature values of the proper motion in the right ascension () and declination () direction from Gaia Collaboration et al. (2018), with a , on areas of identifiable evolutionary features within the Sgr CMD which display the tallest spike of Sgr members.
We then apply a Gaussian fit to stars not identified within our evolutionary features to provide a general cleaning to limit as many Milky Way contaminants within those regions as well, but increasing the two-fold to for a more liberal cleaning. We remove a total of 1,467,943 likely background Milky Way contaminants from the Sgr core, leaving a sample of 515,689 member stars of the Sgr core. Figure 2 visualizes the separation of Sgr-identified members (in red) in the left-most panel and identified MW contamination (in blue).
For M54, there are fewer MW background contaminants visible (in part because our sample is smaller), yet we apply the same cleaning method as before for both the evolutionary subsamples and non-identified evolutionary subsamples. We reuse our prior literature values from Gaia Collaboration et al. (2018) along with the same as there are no prior literature measurements for the proper motion for M54. We visualize the identified Gaussian fits to our histogram evolutionary subsamples in the right-most subplot in Figure 2. We remove a total of 3,340 likely background Milky Way contaminants from M54, leaving a sample of 4,512 member stars of M54.
With our final samples for the Sgr core and M54, we present an overview of the mean astrometric parameters and number of stars within each evolutionary subsample in Table 3. The mean parallax of the Sgr core from Monaco et al. (2004); Minelli et al. (2023) (calculated from the heliocentric distance) was mas. Most of our parallax measurements lie somewhat near this value of mas, however reliable distances cannot be obtained by inverting the parallax beyond kpc, causing the discrepancy in our calculated distance measurements.
4 Red Clump Distance Determination
Red clump stars are the more numerous metal-rich equivalent of horizontal branch stars, found to be standard candles in Paczyński & Stanek (1998), where absolute luminosity weakly depends on age and chemical composition (Udalski, 2000; van Helshoecht & Groenewegen, 2007; Groenewegen, 2008; Girardi, 2016). To support the precision of our membership samples for the Sgr core and M54, we perform a preliminary distance determination to these objects using RC stars as a standard candle.
We cross-match our Sgr core and M54 sample with the Two Micron All Sky Survey Point Source Catalog (2MASS; Cutri et al., 2003; Skrutskie et al., 2006) with a matching 5 arcseconds radius to obtain , , and -band magnitudes of the sources using the TOPCAT software (Taylor, 2005). 2MASS was an all-sky survey in the near-infrared; thus, cross-matching from , which is in the optical, to 2MASS within the NIR allows us to correct for less extinction and reddening as the , , and -bands are less affected by interstellar dust (Skrutskie et al., 2006).
We account for extinction using the dustmaps package (Green, 2018), utilizing the latest version of the three-dimensional dust map Bayestar (Bayestar19; Green et al., 2015, 2018, 2019). Bayestar19 traces dust reddening as a function of angular position and distance along a line of sight to each star. It is probabilistic, reporting how reddening increases or decreases along that sight line (Green et al., 2019). We take the mean value of the dust reddening to our Sgr core and M54 selection stars and convert the given dust reddening to extinction in 2MASS filters using the coefficients reported in Table 1 of Green et al. (2019). We subtract the calculated extinction values from our and magnitudes to correct for extinction.
Following the methodology from Stanek & Garnavich (1998), we create a histogram for both the Sgr core and M54, from our extinction-corrected apparent magnitudes. The peak magnitude within the distribution represents the apparent magnitude of the RC population (). We use the absolute magnitude of the RC in the 2MASS band from Hawkins et al. (2017) of mag to find the distance modulus of each system:
We use a Monte Carlo method to determine the uncertainties on our distance moduli and their corresponding distances. For each distribution of for the Sgr core and M54, we generated a KDE with a bandwidth value of 0.04. From our KDEs, we generated 10,000 resampled distributions over the range our RC was located, ending with a distribution of apparent magnitudes () for our objects. We then modeled the extinction values , generating a distribution based on our results from Bayestar19, resampling 10,000 times. From the absolute magnitude of the RC in the 2MASS band from Hawkins et al. (2017), we generate a normal distribution with a mean of mag and a standard deviation of mag. We calculate the distance moduli for the Sgr core and M54 using these distributions, along with their heliocentric distances, extracting the 16th, 50th, and 84th percentiles to provide the distance moduli to the Sgr core and M54 with errors, along with their distances.
We report a final distance modulus of mag and mag. These distance moduli correspond to a heliocentric distance of kpc for the Sgr Core and kpc for M54.
5 The Connection of the Sgr Core and M54
We isolated the Sgr core and Messier 54 spatially using DR3 astrometry, applied general quality cuts from Gaia Collaboration et al. (2021), and cleaned Milky Way contaminants from our sample by identifying member stars within the histograms of proper motion in right ascension direction and declination direction . We also calibrated distances to both of these systems using a red clump distance determination. Both distance moduli found using red clump distance methods of mag and mag fall well within literature values from Mateo et al. (1995); Monaco et al. (2004); Kunder & Chaboyer (2009); Sollima et al. (2010), as shown in Figure 4, and prove both the precision of our sample and effectiveness of red clump distance methods. Our measured distance moduli to both systems and their respective errors overlap, suggesting that the Sgr core and M54 could occupy the same space. However, as this method only assigns one distance to the entire system–more analysis is needed. A future look into individual star distance determination methods would provide conclusive evidence if M54 is projected onto the area of Sgr’s nucleus.
We corroborate our membership selection using available APOGEE DR17 metallicities and radial velocities from our member stars from the Sgr core and M54. From our Sgr core member selection, we have 1489 stars that have radial velocities and metallicities, and from M54 we have a total of 97 member stars. We show the distribution of radial velocities and metallicities in Figure 5.
We report a range of [Fe/H] of to dex for the Sgr core, with a mean radial velocity of . For M54, we report a [Fe/H] range of to dex, with a mean radial velocity of . Visually, the radial velocities for the Sgr core and M54 are similar, with only difference in their mean. Meanwhile, the Sgr core has a wide-range of [Fe/H], with a high of nearly solar [Fe/H], to a low of dex, matching in the range found in Siegel et al. (2007); Alfaro-Cuello et al. (2019). M54 has a binomial distribution of [Fe/H], matching the findings in Minelli et al. (2023). We find a smaller range of [Fe/H] within M54 with a dex range difference from the Sgr core. We were unable to perform a comparison to the found one dex difference of median [Fe/H] as found in Hayes et al. (2020) did in the streams, due to the slight bimodal distribution featured in both the Sgr core and M54.
We use the line integral convolution (LIC) method, which blurs white noise textures along a vector field to visualize overall flow patterns, of the Sgr core and M54 in the right ascension direction and declination direction . We present the visualization in Figure 6, and notice the same dynamics in both systems–member stars concentrated towards decreasing x-direction, aligning with the pull of the leading arm of Sgr.
The stellar density contours present in the bottom panel of Figure 1 instead showcase a concentration in the central area of the Sgr core near M54, implying decoupling between the motion of the Sgr core its central concentration peak, possibly indicating that M54 remains untouched despite being embedded in the Sgr core.
Evidence supporting the theory that M54 formed independently of the Sgr core comes from their respective red giant branch from their CMDs made from our sample in Figure 3. In the bottom panels, where the RGB and AGB branches are indicated with APOGEE DR17 [Fe/H], the Sgr core displays a rough gradient along its RGB and corresponding AGB branches, marked by three different metallicities. This lines up with the three pericentric interactions with the core of the Milky Way that kick-started star formation in both galaxies (de Boer et al., 2015; Ruiz-Lara et al., 2020). The CMD of M54 lacks these three distinct RGB and AGB branches, instead bearing two distinct branches as found in Figure 2 of Mucciarelli et al. (2017). This suggests that M54 only went through one or two pericentric passages, enriching the cluster, kick-starting star formation, implying M54 infell during the first passage of Sgr. Similarly, the density contours present in the bottom right panel of Figure 1 and the jump from the uniform center to a scattered value outside the first density contour lines with the central density cusp found in Monaco et al. (2005) validate our sample further.
The most recent paper for a Sgr core membership selection was Minelli et al. (2023), where they identified 450 member stars in Early Data Release 3 (EDR3). Cross-matching with our sample, we recover 427 stars from Minelli’s sample. From those 427 stars, none have [Fe/H] measurements from APOGEE DR17, and there are no DR3 [M/H] measurements using General Stellar Parametrizer from Spectroscopy (GSP-Spec) for our sample. There are [M/H] measurements for these 427 within DR3 General Stellar Parametrizer from Photometry (GSP-Phot), however these values exhibit substantial biases compared to literature values and are only useful at a qualitative level especially for red giant branch stars with low quality parallax measurements (Andrae et al., 2023). Other attempts to gain [Fe/H] measurements from The Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) and The GALactic Archaeology with HERMES (GALAH) fail to capture Sgr dSph core and M54 member stars within our sample, unfortunately not allowing us a direct comparison of Minelli et al., 2023 [Fe/H] measurements taken from the Fibre Large Array Multi Element Spectrograph (FLAMES) on the Very Large Telescope (VLT).
6 Conclusion
In this paper, we have used DR3 and APOGEE DR17 to produce the largest catalog of stars from the Sgr Core and Messier 54, separating member stars from the heavily contaminated area from foreground and background Milky Way stars. We present the precision of our samples using red clump distance determinations using magnitudes from 2MASS. Our main conclusions are as follows:
We demonstrate the strength of DR3 astrometry combined with APOGEE DR17 spectra to separate the Sgr core with M54, along with a multi-Gaussian fit to reduce MW contaminants for systems near the disk. We have a membership catalog of 515689 stars in the Sgr core and 4512 in M54.
We measure the distance to both the Sgr core and M54 using red clump distancing methods from a cross-matched sample with 2MASS, finding a heliocentric distance of kpc for the Sgr Core and kpc for M54. Our distance measurements agree with the literature but are more precise due to the larger member sample identified in our work.
We analyze the motion, density contours, and [Fe/H] of the Sgr core and M54 to understand whether M54 formed independently and was captured during a passage, or is the Sgr nucleus. Preliminary distance measurements and the dynamics of our sample imply they co-exist, but the chemical abundances suggest the systems formed independently.
The results presented here provide the first glimpse into the full dynamics of the Sgr core and its relationship to M54 in DR3. Utilizing this large membership sample and cross-matching with current and future all-sky ground-based and space-based surveys will allow further study into the dynamics of the Sgr core and M54. Combined with masses and ages from asteroseismic data sets, we can further understand the formation, and relationship of the Sgr core and M54.
7 Acknowledgments
E.K.H.T.T. acknowledges support from the Research Experience for Undergraduates program at the Institute for Astronomy, University of Hawai‘i-Mānoa funded through the National Science Foundation grant 6109694. E.K.H.T.T. also acknowledges support from the Whitman Internship Grant through Whitman College. E.K.H.T.T. would like to thank the Institute for Astronomy for their hospitality during this project.
E.K.H.T.T. also gratefully acknowledges support for this research from the National Science Foundation’s Research Experience for Undergraduates program PHY-2244258 at Texas Christian University.
E.K.H.T.T. would like to additionally extend gratitude to Dr. Robyn Sanderson from the University of Pennsylvania, Dr. Sukanya Chakrabarti from the University of Alabama, Huntsville, Dr. Marc Hon from the Massachusetts Institute of Technology, Dr. Andrea Dobson from Whitman College, Dr. Matthew Shetrone from the University of Californa, Santa Cruz, and Aldo Sepuvelda from the University of Texas at Austin School of Law for their insightful conversations and feedback on this work.
D.H. acknowledges support from the Alfred P. Sloan Foundation and the National Science Foundation (AST-2009828).
P.M.F. acknowledges support from the National Science Foundation (AST-2206541).
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions.
SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org.
SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics — Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
8 Appendix
8.1 Gaia DR3 ADQL Queries
The following DR AQDL queries produce our spatial sample for the Sgr dSph core and Milky Way contamination, respectively, in full.
SELECT TOP 50000000
gaia_source.designation, gaia_source.source_id, gaia_source.ra, gaia_source.ra_error,
gaia_source.dec, gaia_source.dec_error, gaia_source.parallax, gaia_source.parallax_error,
gaia_source.parallax_over_error, gaia_source.pm, gaia_source.pmra, gaia_source.pmra_error,
gaia_source.pmdec, gaia_source.pmdec_error, gaia_source.visibility_periods_used,
gaia_source.phot_g_mean_mag, gaia_source.bp_rp, gaia_source.radial_velocity,
gaia_source.radial_velocity_error, gaia_source.rv_template_fe_h,
gaia_source.phot_variable_flag, gaia_source.l, gaia_source.b, gaia_source.non_single_star,
gaia_source.has_epoch_rv, gaia_source.teff_gspphot, gaia_source.logg_gspphot,
gaia_source.mh_gspphot, gaia_source.distance_gspphot, gaia_source.ag_gspphot,
gaia_source.ebpminrp_gspphot
FROM gaiadr3.gaia_source
WHERE CONTAINS(
POINT(’ICRS’, gaiadr3.gaia_source.ra, gaiadr3.gaia_source.dec),
CIRCLE(’ICRS’,
COORD1(EPOCH_PROP_POS(283.8292,-30.5453,0,-2.6500,-.8800,140.0000,2000,2016.0)),
COORD2(EPOCH_PROP_POS(283.8292,-30.5453,0,-2.6500,-.8800,140.0000,2000,2016.0)), 4)
)=1
AND gaia_source.phot_g_mean_mag IS NOT NULL
AND gaia_source.parallax IS NOT NULL
AND gaia_source.bp_rp IS NOT NULL
SELECT TOP 50000000 gaia_source.designation, gaia_source.source_id, gaia_source.ra, gaia_source.ra_error, gaia_source.dec, gaia_source.dec_error, gaia_source.parallax, gaia_source.parallax_error, gaia_source.parallax_over_error, gaia_source.pm, gaia_source.pmra, gaia_source.pmra_error, gaia_source.pmdec, gaia_source.pmdec_error, gaia_source.visibility_periods_used, gaia_source.phot_g_mean_mag, gaia_source.bp_rp, gaia_source.radial_velocity, gaia_source.radial_velocity_error, gaia_source.rv_template_fe_h, gaia_source.phot_variable_flag, gaia_source.l, gaia_source.b, gaia_source.non_single_star, gaia_source.has_epoch_rv, gaia_source.teff_gspphot, gaia_source.logg_gspphot, gaia_source.mh_gspphot, gaia_source.distance_gspphot, gaia_source.ag_gspphot, gaia_source.ebpminrp_gspphot FROM gaiadr3.gaia_source WHERE CONTAINS( POINT(’ICRS’, gaiadr3.gaia_source.ra, gaiadr3.gaia_source.dec), CIRCLE(’ICRS’, 259.65, -18.47, 3))=1 AND gaia_source.phot_g_mean_mag IS NOT NULL AND gaia_source.parallax IS NOT NULL AND gaia_source.bp_rp IS NOT NULL
8.2 Flattened Coordinates
Below we showcase the equations to compute the flattened coordinates for the Sgr core and account of the elliptical shape of the core.
8.3 Candidate Members Table
Table 4 and Table 5 (full tables available at CDS), contain the data for identified members of the Sgr core and Messier 54. Below we show an excerpt of the first five rows from each system.
| Gaia Source_id | ||||||
| 6736253232769633408 | 282.81 | -31.56 | -2.41 | -1.36 | 17.96 | 1.24 |
| 6736253542006710528 | 282.80 | -31.54 | -2.91 | -1.07 | 17.92 | 1.37 |
| 6736253885604674688 | 282.72 | -31.54 | -2.50 | -1.43 | 17.82 | 1.29 |
| 6736254126121627648 | 282.68 | -31.53 | -2.57 | -1.12 | 17.97 | 1.29 |
| 6736254233517136128 | 282.75 | -31.54 | -0.57 | -2.72 | 17.91 | 1.12 |
Full table is available at the CDS.
| Gaia DR3 Source_id | ||||||
| 6760412629965372928 | 283.83 | -30.56 | -2.89 | -1.52 | 17.97 | 1.40 |
| 6760412629966580096 | 283.84 | -30.56 | -2.70 | -1.08 | 17.84 | 1.26 |
| 6760412629969429632 | 283.83 | -30.57 | -2.79 | -1.07 | 17.94 | 1.29 |
| 6760412634271816448 | 283.84 | -30.57 | -2.65 | -1.70 | 17.96 | 1.30 |
| 6760412634289471488 | 283.84 | -30.57 | -2.56 | -1.40 | 18.01 | 1.40 |
Full table is available at the CDS.